how much salt must be added to 800 grams of a solution that is 15% to obtian a solution that is 20% salt?
Answer by mananth(16946)

how much salt must be added to 800 grams of a solution that is 15% to obtian a solution that is 20% salt?
let x grams of salt be added.
x+0.15*800=(x+800)*0.2
x+120=0.2x+160
0.8x=40
x=40/0.8
50 grams